OpenStudy (anonymous):

If x3 + y3 = 9 and dx/dt = -3, then what is dy/dt when x = 1 and y = 2?

OpenStudy (anonymous):

@RyanRyan What is the dx/dt there for?

OpenStudy (anonymous):

Regardless, @satellite73 has pretty much showed you exactly what to do.

OpenStudy (anonymous):

actually i was completely wrong!!

OpenStudy (anonymous):

no idea, it threw me off

OpenStudy (anonymous):

1. Differentiate both sides with respect to x 2. Plug in x = 1 and y = 2. 3. Solve for y'. Done.

OpenStudy (anonymous):

Wait what? is there more to this problem? lol

OpenStudy (anonymous):

Oh wait I get it. x and y are functions of t. Lol I forgot about that.

OpenStudy (anonymous):

no no i made a mistake by not reading properly you are supposed to assume that BOTH \(x\) and \(y\) are some unknown functions of \(t\)

OpenStudy (anonymous):

\[3x^2x'+3y^2y'=0\] is the start

OpenStudy (anonymous):

thats the exact problem.. It is extra credit so maybe she is making it confusing? or its just late, I dont know, this particular problem is multi choice too

OpenStudy (anonymous):

then pout \(x'=-3,x=1,y=2\) and solve for \(y'\)

OpenStudy (anonymous):

no problem is perfectly normal you are not asked for \(\frac{dy}{dx}\) but rather \(\frac{dy}{dt}\)

OpenStudy (anonymous):

Differentiate both sides with respect to t:\[\bf 3x^2\frac{ dx }{ dt }+3y^2\frac{ dy }{ dt }=0\]Now plug in: dx/dt = -3, x = 1, y = 2:\[\bf 3(1)^2(-3)+3(2)^2\frac{ dy }{ dt }=0\]Now rearrange and solve for dy/dt.

OpenStudy (anonymous):

@RyanRyan

OpenStudy (anonymous):

@RyanRyan You see what I did there? Do you have any questions/trouble understanding?

OpenStudy (anonymous):

yea, 3/4.. thanks guys

OpenStudy (anonymous):

@RyanRyan Good job. Your answer is correct =D.